OpenStudy (anonymous):

-8x+8y=-16 when x is 0

OpenStudy (mathlegend):

@VictoriaJustMeGriffin plug it in then... -8(0) + 8y = -16

OpenStudy (mathlegend):

Now, what can we do next?

OpenStudy (anonymous):

umm we do -8 x 0 which give u 0 so I dnt kno wats next

OpenStudy (mathlegend):

Which means it is gone.. so now we have.. 8y = -16

OpenStudy (mathlegend):

So can't we divide now?

OpenStudy (mathlegend):

Solve for y.

OpenStudy (anonymous):

umm

OpenStudy (mathlegend):

It is being multiplied so we must divide to get y by itself...

OpenStudy (mathlegend):

\[\frac{ 8y }{ 8 } = \frac{ -16 }{ 8 }\]

OpenStudy (mathlegend):

Now, solve for y.

OpenStudy (anonymous):

-2

OpenStudy (mathlegend):

Good.

OpenStudy (anonymous):

thanx :)